OpenStudy (anonymous):

Verify the identity. tan(x + (π/2)) = -cot x

OpenStudy (anonymous):

Use the Tan (A+B) Formula

OpenStudy (anonymous):

it wouldn't be the cofunction identities?

OpenStudy (jdoe0001):

I don't think is one , though I gather it should be, but who knows, if it equates, then yeah it should be :)

OpenStudy (anonymous):

well i mean it is a reduction identity, but i dont know how to prove it

OpenStudy (jdoe0001):

well, use the tan(a+b) as FutureMathProfessor suggested

OpenStudy (anonymous):

i did and got (tan x + tan (π/2))/ (1 - tanxtanπ/2)

OpenStudy (jdoe0001):

... I expanded by sine/cosine and got ... nothing :/

OpenStudy (jdoe0001):

wait. I just gimme a sec

OpenStudy (jdoe0001):

$$ tan\pmatrix{x+\cfrac{\pi}{2}} \implies \cfrac{sin\pmatrix{x+\cfrac{\pi}{2}}}{cos\pmatrix{x+\cfrac{\pi}{2}}} \implies \cfrac{cos(x)}{-sin(x)} $$

OpenStudy (jdoe0001):

and those 2 are identities :)

OpenStudy (anonymous):

ok thankyou